OpenStudy (anonymous):

HELP PLEASE :( i've been working on this for almost the whole night and i'm still stumped :( An urn contains five red and three blue chips. Suppose that four of these chips are selected at random and transferred to a second urn, which was originally empty. If a random chip from this second urn is blue, what is the probability that two red and two blue chips were transferred from the first urn to the second urn?

OpenStudy (anonymous):

@.Sam. @ash2326 @ParthKohli @TuringTest @apoorvk @Carl_Pham @Coolsector @estudier @hartnn @helder_edwin @Hero @matricked @satellite73 @sauravshakya @Zekarias help please :(

OpenStudy (anonymous):

i'm reviewing for my upcoming midterms and i got this from a past exam... it's scaring me :( i'm familiar with conditional probability and i could answer majority of the questions in my book just fine, but i find this too complicated. :(

OpenStudy (ash2326):

We have to use Bayes' theorem here. Have you studied that?

OpenStudy (anonymous):

oh! really? we weren't taught that. maybe there's a way of solving it without the need of that? like a tree diagram? or is it impossible without that theorem?

OpenStudy (ash2326):

Umm, I don't know any other way It's conditional probability, refer this first. Then we'll work on the problem http://en.wikipedia.org/wiki/Bayes'_theorem

OpenStudy (anonymous):

oh! i'm aware of the simple form of the bayes' theorem! i wasn't aware it was called that. i actually tried to use it in this, but i had a hard time identifying the components:(

OpenStudy (ash2326):

Ok, let's work :)

OpenStudy (anonymous):

i tried to solve for P(2 red and 2 blue chips were transferred from the first urn to the 2nd urn | random chip from second urn is blue) but i just got so lost :( i really appreciate your patience in helping me through this @ash2326 !!!! thank you already!!

OpenStudy (ash2326):

We need to find probability of 2 R and 2 Blue being transferred from urn A to B, provided random chip is blue. Possible combinations when 4 chips are transferred from A to B. Let the combinations be 2R and 2B= X 1R and 3B= Y 3R and 1B=Z 4R=U Do you get this?

OpenStudy (ash2326):

I'm here to help, you're welcome.

OpenStudy (ash2326):

I've just given aliases to the combinations

OpenStudy (ash2326):

@dydlf are you here?

OpenStudy (anonymous):

yes i am! yes i get you so far.

OpenStudy (ash2326):

Ok, let's continue. Feel free to interrupt where you have doubt

OpenStudy (anonymous):

sure!

OpenStudy (ash2326):

So we need to find \[P(X/B)=\text{ probability of 2 R and 2B being transferred when a random chip is blue}\] \[\small {P(X/B)=\frac{P(B/X)\times P(X)}{P(B/X)\times P(X)+P(B/Y)\times P(Y)+P(B/Z)\times P(Z)+P(B/U)\times P(U)}}\] Do you get this? @dydlf

OpenStudy (anonymous):

hmmm let me refer to the wiki page you sent. wait

OpenStudy (ash2326):

sure

OpenStudy (anonymous):

sorry, lost connection. but okay! i see what you mean now!

OpenStudy (ash2326):

Now let's find the individual probabilities \[P(B/X)\times P(X)\] Can you tell me the value of this?

OpenStudy (ash2326):

or How do we find it?

OpenStudy (anonymous):

hmm doesn't P(B|X)×P(X) mean the intersection of B and X? when i tried solving this, that's where i got stuck since i didn't know how to get this. hmmm P(X) is 5C2×3C2/10C4 ? P(B|X) is like P(B intersection X)/P(X) right? hmmm 2C1/(5C2×3C2/10C4) ?

OpenStudy (ash2326):

Let's work step by step probability of getting blue when 2 blue and 2 red are transferred =P(B/X) It's just the probability of getting one blue, when there are 2 red and 2 blue chips \[P(B/X)=\frac {2}{4}\]

OpenStudy (anonymous):

gosh i meant 5C2×3C2/8C4 not 5C2×3C2/10C4

OpenStudy (ash2326):

Now we'll find P(X), that's choosing 2 red and 2 blue from 5 red and 3 blue \[P(X)=\frac{^5C_2 \times ^3C_2}{^8C_4}\] Yeah you got P(X) right

OpenStudy (ash2326):

@dydlf Do you get how I found P(B/X) ?

OpenStudy (ash2326):

Take your time, I'm here :)

OpenStudy (anonymous):

hmm yup! but i'm just wondering, what's the diff of let's say P(B|X) and P(B intersection X)? like in words?

OpenStudy (ash2326):

P(B intersection X) is probability of occurrence of B and X event simultaneously P(B/X) is the conditional probability of occurrence of B when X has already occurred

OpenStudy (ash2326):

Infact P(B/X)=P( B inter X)/P(X)

OpenStudy (anonymous):

yes i understand that, but i don't really understand the concept like for solving P(B|X) earlier, if you noticed, i used that equation you just mentioned up there, but i had a hard time looking for P(B inter X) since i think i don't really understand the concept

OpenStudy (anonymous):

i know it has to be in both B and X but... gosh i just got confused there.

OpenStudy (ash2326):

We don't need to find P( B inter X) here Finding P(B/X ) is straightforward . See how I found it above

OpenStudy (phi):

Here is one way to do it without explicitly invoking Bayes Theorem

OpenStudy (anonymous):

oh, i was just wondering. because i usually mix the 2 up....

OpenStudy (ash2326):

Do you get till here?

OpenStudy (phi):

Given that my solution does not match Bayes Thm approach, I conclude there is a flaw in my argument, and you really do have to grind through the problem.

OpenStudy (anonymous):

Is it 6/13 ??? The answer... I'm getting that only... @dydlf

OpenStudy (phi):

No, the answer, if you finish up Ash's approach, is 4/7 the numerator is (2/4)* (30/70) the denominator is (2/4)*(30/70) + (3/4)*(5/70) + (1/4)*(30/70) + 0*(5/70)

OpenStudy (phi):

I believe it comes down to how to interpret "If a random chip from this second urn is blue..." If this means "A chip selected from the 2nd urn is blue..." then we get 4/7 If we interpret it as "At least one chip in the 2nd urn is blue..." we get 6/13

OpenStudy (anonymous):

I don't think... my approach makes it 6/13 only... maybe some flaw is there in his or mine... I think he doesn't need to consider the case when 4 red balls are chosen... because it is mentioned in the question that... *a random chip from this second urn is blue* which makes me think there's at least one blue ball in the second urn...

OpenStudy (anonymous):

*chip*

OpenStudy (phi):

Yes, I agree it is a subtle question. This is a question on an old exam, and I would like to see its solution write-up.

OpenStudy (anonymous):

Even if I consider your 1st approach... I get 3/7 as against yours 4/7...

OpenStudy (phi):

How ?

OpenStudy (anonymous):

 2R, 2B ===> 5C2* 3C2 = 30 3R, 1B ===> 5C3*3C1 = 30 1R, 3B ===> 5C1*3C3 = 5 4 Red ===> 5C4 = 5 P(2R, 2B) = 30/ (30+30+5+5) = 3/7 Please check if there's some flaw...